In aggregate planning, the forecasted demands for the planning horizon are 1280, 640, 900, 1200, and 1400. Currently, there are 300 units of inventory on hand at the beginning of the 1st period and it is required that an ending inventory of 500 is needed by the end of the last period. After absorbing the starting and ending inventories, the actual production volume needed for the 1st period is ____, for the last period is _____.
Answer #1

Given there are 5 planning horizons.

Forecasted demands for the planning horizon are :-

Planning Horizon                  Forecasted demands

            1                                          1280

            2                                           640

            3                                           900

            4                                          1200

            5                                          1400

Inventory on hand at the beginning of the 1st period = 300

Ending inventory required by the end of 5th period = 500

Hence the actual production volume for 1st period = (1280 - 300) = 980

The actual production volume for 5th period = (1400 + 500) = 1900